# Pumping and Steady Streaming driven by Two-Frequency Oscillations of a Cylinder

### v1

@@
The classical problem of steady streaming induced by an oscillating object has been studied extensively, but prior work has focused almost exclusively on single-frequency oscillations, which result in symmetric, quadrupole-like flows. Here we demonstrate that dual-frequency oscillations induce asymmetric steady streaming with a non-zero net flux in a direction determined by the polarity of the oscillation – the oscillator serves as a pump. We use numerical simulations and asymptotic analysis at low Reynolds number to examine 2D steady streaming around a cylinder, first focusing on frequency ratio two. The computational experiments show asymmetrical streaming and pumping, i.e., net flux downstream. It is well known from asymptotic analysis that steady streaming is second order in amplitude, and we show pumping occurs at third order. We then extend the analysis to general frequency ratios, where we give necessary conditions for pumping and predict the order in amplitude at which pumping occurs. Finally, we corroborate the theoretical results with computational simulations for different frequency ratios, and we discuss the implications for using dual-mode vibrations to pump fluids in lab-on-a-chip and other applications.
@@ 1. Introduction
Steady streaming induced by an oscillating object is a classical phenomenon that has been studied analytically and experimentally (Riley, 2001). The experimental study of steady streaming around a circular cylinder began with Carrière (1929) and Andrade (1931), who examined periodically oscillating air around a cylinder and reported the now familiar quadrupole-like flow. Schlichting (1932) carried out the first asymptotic analysis of steady streaming around a cylinder, and he compared the solution with his experiments on a vibrating cylinder in water. In the small amplitude limit, he matched the inner boundary layer with the outer potential flow, and his solution became the basis for future analysis. For example, Wang (1968) adapted the layer analysis to include the curvature of the boundary, which is relevant at higher order. Riley (1965) and Stuart (1966) analyzed the problem at high Reynolds number where a second boundary layer exists, and their results agreed well with Schlichting’s experiments. Finally, Holtsmark et al. (1954) analyzed the problem at small Reynolds number, where a boundary layer is absent, and solved the flow everywhere as a regular perturbation problem.
@@
Though the problem of steady streaming has been studied extensively, past works have primarily focused on single-frequency oscillations, with comparatively less attention to multi-frequency oscillations (Davidson & Riley, 1972; Kotas et al., 2008). Davidson & Riley (1972) analyzed steady streaming of multi-frequency oscillation around a cylinder and found the resultant flow is the superposition of steady streaming flows induced by each single-frequency oscillation. Similarly, Kotas et al. (2008) experimentally examined the steady flow around a sphere made to oscillate with two frequencies, and they found that the observed flows were the sum of the streaming flows from the respective single-frequency oscillations. The experiments in both of these works were carried out for small amplitude motion in which the streaming velocity is small, and the asymptotic analysis was only performed up to second order in amplitude. There are reasons to believe that imposition of multiple oscillatory frequencies might break the spatial symmetry and induce net flow in one direction at higher order in amplitude.
@@
This work is inspired by recent experiments on a different physical system: an object sliding on a surface, undergoing two-frequency lateral oscillation (Hashemi et al., 2022). Specifically, the surface displacement was $(\ell/2)[\sin(\omega t)+\sin(\alpha\omega t)]$, where $\ell$ is the amplitude, $\omega$ is the frequency, and $\alpha$ is a ratio of the two frequencies. For particular $\alpha$-values, the object exhibited a net translation. To help understand the experimental observations, they analyzed a model in which the only force considered is the surface contact described by Coulomb’s friction law. Representing the frequency ratio $\alpha$ as $p/q$, they derived a necessary condition for net motion: one of $p$ or $q$ is odd and the other is even, and validated this prediction with experiments. For example, $\alpha=2$, $3/2$, and $1/2$ showed net motion while $\alpha=1$, $3$, and $5/3$ did not. We note the same necessary conditions for directed motion were derived earlier by Reznik & Canny (2001) using a simpler friction model (Reznik et al., 1997). Additional analysis and experiments for oscillations with different phase and amplitude were carried out in Zhang et al. (2024a), and Hui et al. (2024) further corroborated the theory with experiments with granular media. Asymptotic analysis of a related problem showed that the translation speed of the object is third order in amplitude (Hashemi et al., 2024), which motivates us to reexamine multi-frequency steady streaming to consider effects beyond second order.
@@ 2.1. Cylinder Motion
We use two different equivalent formulations of the problem: one is in which the cylinder moves and the other in which the cylinder is stationary and the far field flow oscillates. For the moving cylinder formulation, a cylinder of radius $R$ and center $(X(t),0)$ moves in a 2D viscous fluid. The horizontal position of the cylinder’s center is prescribed as
@@
For analysis is it convenient to consider the cylinder fixed at the origin and impose a the horizontal flow at at infinity of strength
@@ 2.2. Nondimensionalization
The Reynolds number, Re $R^{2}\Omega/\nu$, is based on the length scale of the cylinder’s radius and corresponds to the boundary layer in which viscous forces remain relevant. The streaming Reynolds number, $\text{Re}_{s}$, is based on length scale of the oscillation amplitude and it characterizes the Reynolds number associated with the streaming flow. The two Reynolds numbers are related by
@@
The flow structure depends on both Re and $\text{Re}_{s}$. We refer to Figure 1 from Chong et al. (2013) (which is adapted from Figure 1 of Wang (1968)) to illustrate the different flow regimes and associated analyses. Briefly, for high Reynolds numbers the flow exhibits a boundary layer (Schlichting, 1932; Wang, 1968), and at high streaming Reynolds numbers, the streaming flow itself develops a second boundary layer (Stuart, 1966; Riley, 1965). At low Reynolds number, there are no boundary layers, and the problem can be analyzed as a regular perturbation (Holtsmark et al., 1954). Though as noted in Chong et al. (2013), the regular perturbation solution from Holtsmark et al. (1954) contains the single boundary layer solutions from Schlichting (1932) and Wang (1968). In this paper, we focus on the low Reynolds number regime. The regular perturbation analysis allows us to compute the structure of solutions beyond second order needed to understand pumping from two-frequency oscillations.
@@ 3.1. Comparing Flow Patterns
We place the cylinder in a square 8 by 8 domain with periodic boundary conditions and solve for the flow at Re =10=10. We solve for 250 periods which is sufficiently long to allow the initial transient to decay, and the resulting flows are visualized by the positions of passive marker particles from the previous 10 periods of time. As expected for the single-frequency case in Figure 1(a), the flow resembles the classical, four vortices associated with steady streaming. However, for two-frequency motion, shown in Figures 1(b,c), the vortices are no longer symmetric and there is a net horizontal flow. The bottom panels show the cylinder horizontal position for each case. In Figure 1(b), the net flow is left-to-right – away from the cylinder, the green particles (most current period) are to the right of the blue particles (past period), indicating a net rightward flow. See movie 1 in the supplementary materials to compare the flow patterns of single-frequency, two-frequency, and reverse two-frequency oscillations.
@@ 3.2. Pumping in a Channel
The results in 1(b) suggest that cylinders oscillating with two frequencies can be used to pump fluids. To test that idea more directly, we repeated the computational experiments in a channel of height 8 and length 32 with no-slip boundary conditions on the top and bottom and periodic conditions in the horizontal direction. Figure 2 is a picture of movie 2 at time $T=250$, and they show the contrast of single and two-frequency motions inside a channel. Again, for the single-frequency case, the four vortices align symmetrically around the cylinder, and particles away from the cylinder remain stationary. However, when the cylinder oscillates with two-frequencies, the vortices lose their symmetry, and fluid is pumped to the right. Applying the negative waveform again reversed the direction of motion (results not shown).
@@
Figure 3(a) shows the flux through the channel,
@@
over $5$ periods for both the single and two-frequency cases for amplitude $\epsilon=0.9$. The flux for the two-frequency case shows a nonzero average flux on the scale of about 10% of the amplitude of the oscillation. We quantify pumping by the time average of the flux over the period $T$:
$\langle Q(t)\rangle\;=\frac{1}{T}\int_{t-T}^{t}F(\tau)\;d\tau.$
@@
Figure 3(c) shows that the steady flux grows like amplitude cubed. For single-frequency oscillations the magnitude of the steady flow is second order in amplitude and the oscillatory component is first order (Holtsmark et al., 1954; Wang, 1968; Chong et al., 2013). In Figure 3(d) we show that for the two-frequency case these same scalings for the size of the steady and oscillatory flows result as measured by the $2$-norm of the period-averaged flow and the time average of the $2$-norm of the velocity, respectively. Previously analysis on steady streaming has only been carried out to second order in amplitude, and so to derive the terms responsible for pumping, we need to extend the analysis to third order.
Figure 3. (a) Flux vs. time is presented for both the single and two-frequency case for amplitude 0.9 and Reynolds number 10. (b) Time-averaged fluxes are shown for varying amplitudes $(\epsilon$). (c) The time-averaged flux is third order in amplitude for two-frequency oscillations. (d) The size of the oscillatory component, $\langle\|u\|_{2}\rangle$, is first order in amplitude, and the size of the steady streaming flow, $\|\langle u\rangle\|_{2}$, is second order in amplitude for the two-frequency oscillation.
@@ 4. Low Amplitude Analysis
We restrict the analysis to low Reynolds numbers because there is no boundary layer and the problem can be analyzed as a regular perturbation. The single-frequency solution has been previously computed through second order in amplitude (Holtsmark et al., 1954; Chong et al., 2013), but extending the analysis to third order is not analytically tractable. However, we derive the structure of the higher order solutions, and, from the structure alone, we can understand why pumping is a third order effect for the two-frequency case.
@@ 4.1. Governing Equations
We solve the problem in the reference frame in which the cylinder is stationary cylinder with an oscillatory flow at infinity. The boundary boundary conditions in terms of the the stream function are
@@ 4.2.1. Single-Frequency Oscillation
$\psi_{1}=\Re\left(a_{1}(r)e^{-it}\right)\sin(\theta).$ (22)
We substitute (22) into the quadratic nonlinearity on the right side of Equation (16) to determine the form of the solution at second order. This computation involves all products of $e^{-it}$ and $e^{it}$, which results in a steady term and terms proportional to $e^{\pm i2t}$. Therefore, at second order, we obtain
@@ 4.2.2. Two-Frequency Oscillation
We express the solution in this second form with three pairs of terms grouped together to highlight the origin of the different terms in the sum. The first two pairs involving a steady term and unsteady term correspond to the single-frequency solutions from each of the two driving frequencies, and the last pair of unsteady terms arises from the interaction between the driving frequencies. Thus as has been reported previously, the steady flow at second order is the superposition of two streaming fields caused by the individual, single-frequency oscillations (Davidson & Riley, 1972; Kotas et al., 2008).
@@
Unlike the single-frequency solution (24), there is a steady from at third order. In the following sections, we explain how the third-order, steady terms produce the observed left-right asymmetry and are responsible for producing a net force that pumps the fluid.
@@ 4.3. Symmetry of Steady Flow
To illustrate how the steady flows at 3rd order create asymmetry, consider the steady solution,
@@
We show that the familiar second order flow has a different spatial symmetry than the third order flow that arise from the interactions of the flows generated by the two different driving frequencies. Specifically, the spatial symmetry arises from the $\theta$-dependence of the stream functions, and, to illustrate, we examine streamlines of the three stream functions,
@@
This choice of $f(r)$ satisfies the boundary conditions on the cylinder and decays to zero at infinity.

### v2

@@
The classical problem of steady streaming induced by an oscillating object has been studied extensively, but prior work has focused almost exclusively on single-frequency oscillations, which result in symmetric, quadrupole-like flows. Here we demonstrate that dual-frequency oscillations induce asymmetric steady streaming with a non-zero net flux in a direction determined by the polarity of the oscillation – the oscillator serves as a pump. We use numerical simulations and asymptotic analysis at small amplitude to examine 2D steady streaming around a cylinder, first focusing on frequency ratio two. The computational experiments show asymmetrical streaming and pumping, i.e., net flux downstream. It is well known from asymptotic analysis that steady streaming is second order in amplitude, and we show pumping occurs at third order. We then extend the analysis to general frequency ratios, where we give necessary conditions for pumping and predict the order in amplitude at which pumping occurs. Finally, we corroborate the theoretical results with computational simulations for different frequency ratios, and we discuss the implications for using dual-mode vibrations to pump fluids in lab-on-a-chip and other applications.
@@ 1. Introduction
Steady streaming induced by an oscillating object is a classical phenomenon that has been studied analytically and experimentally (Riley, 2001). The experimental study of steady streaming around a circular cylinder began with Carrière (1929) and Andrade (1931), who examined periodically oscillating air around a cylinder and reported the now familiar quadrupole-like flow. Schlichting (1932) carried out the first asymptotic analysis of steady streaming around a cylinder, and he compared the solution with his experiments on a vibrating cylinder in water. In the small amplitude limit, he matched the inner boundary layer with the outer potential flow, and his solution became the basis for future analysis. For example, Wang (1968) adapted the layer analysis to include the curvature of the boundary, which is relevant at higher order. Riley (1965) and Stuart (1966) analyzed the problem at high Reynolds number where a second boundary layer exists, and their results agreed well with Schlichting’s experiments. Finally, Holtsmark et al. (1954) analyzed and solved the flow everywhere as a regular perturbation problem.
@@
Though the problem of steady streaming has been studied extensively, past works have primarily focused on single-frequency oscillations, with comparatively less attention to multi-frequency oscillations (Davidson and Riley, 1972; Kotas et al., 2008). Davidson and Riley (1972) analyzed steady streaming of multi-frequency oscillation around a cylinder and found the resultant flow is the superposition of steady streaming flows induced by each single-frequency oscillation. Similarly, Kotas et al. (2008) experimentally examined the steady flow around a sphere made to oscillate with two frequencies, and they found that the observed flows were the sum of the streaming flows from the respective single-frequency oscillations. The experiments in both of these works were carried out for small amplitude motion in which the streaming velocity is small, and the asymptotic analysis was only performed up to second order in amplitude. There are reasons to believe that the imposition of multiple oscillatory frequencies might break the spatial symmetry and induce net flow in one direction at higher order in amplitude.
@@
This work is inspired by recent experiments on a different physical system: an object sliding on a surface, undergoing two-frequency lateral oscillation (Hashemi et al., 2022). Specifically, the surface displacement was $(\ell/2)[\sin(\omega t)+\sin(\alpha\omega t)]$, where $\ell$ is the amplitude, $\omega$ is the frequency, and $\alpha$ is a ratio of the two frequencies. For particular $\alpha$-values, the object exhibited a net translation. To help understand the experimental observations, they analyzed a model in which the only force considered is the surface contact described by Coulomb’s friction law. Representing the frequency ratio $\alpha$ as $p/q$, they derived a necessary condition for net motion: one of $p$ or $q$ is odd and the other is even, and validated this prediction with experiments. For example, $\alpha=2$, $3/2$, and $1/2$ showed net motion while $\alpha=1$, $3$, and $5/3$ did not. We note the same necessary conditions for directed motion were derived earlier by Reznik and Canny (2001) using a simpler friction model (Reznik et al., 1997). Additional analysis and experiments for oscillations with different phase and amplitude were carried out in Zhang et al. (2024b), and Hui et al. (2024) further corroborated the theory with experiments with granular media.
@@ 2.1. Cylinder Motion
We use two related formulations of the problem: one is in which the cylinder moves and the other in which the cylinder is stationary and the far-field flow oscillates. For the moving cylinder formulation, a cylinder of radius $R$ and center $(X(t),0)$ moves in a 2D viscous fluid. The horizontal position of the cylinder’s center is prescribed as
@@
For analysis it is convenient to consider the cylinder fixed at the origin and impose a horizontal flow at infinity of strength
@@
The two formulations are related by a change in reference frame. Switching our frame of reference from the inertial frame where the cylinder is oscillating to the non-inertial frame in which the cylinder is fixed results in a fictitious body force (Batchelor, 2000),
$F=-\rho U^{\prime}(t)\boldsymbol{e_{x}},$ (4)
where $\rho$ is the mass density and $\boldsymbol{e_{x}}$ is the direction vector in the horizontal direction. In the non-inertial frame, we incorporate the fictitious body force (4) in the modified pressure defined as,
$\tilde{p}=p+\rho xU^{\prime}(t).$
For simplicity, we denote the modified pressure as $p$ below.
@@ 2.2. Nondimensionalization
The Reynolds number, Re $R^{2}\Omega/\nu$, is based on the length scale of the cylinder’s radius and corresponds to the boundary layer in which viscous forces remain relevant. The streaming Reynolds number, $\text{Re}_{s}$, is based on the length scale of the oscillation amplitude and it characterizes the Reynolds number associated with the streaming flow. The two Reynolds numbers are related by
@@
The flow structure depends on both Re and $\text{Re}_{s}$. We refer to Figure 1 from Chong et al. (2013) (which is adapted from Figure 1 of Wang (1968)) to illustrate the different flow regimes and associated analyses. Briefly, for high Reynolds numbers the flow exhibits a boundary layer (Schlichting, 1932; Wang, 1968), and at high streaming Reynolds numbers, the streaming flow itself develops a second boundary layer (Stuart, 1966; Riley, 1965). At low Reynolds number, there are no boundary layers, and the problem can be analyzed as a regular perturbation (Holtsmark et al., 1954). Though as noted in Chong et al. (2013), the regular perturbation solution from Holtsmark et al. (1954) contains the single boundary layer solutions from Schlichting (1932) and Wang (1968), and is thus appropriate provided that the streaming Reynolds number is small. In this paper, we focus on the low streaming Reynolds number regime $(\text{Re}_{s}\ll 1)$ so that the problem can be analyzed as a regular perturbation, which allows us to compute the structure of solutions beyond second order needed to understand pumping from two-frequency oscillations.
@@ 3.1. Comparing Flow Patterns
We place the cylinder in a square 8 by 8 domain with periodic boundary conditions and solve for the flow at Re =10=10. We solve for 250 periods which is sufficiently long to allow the initial transient to decay, and the resulting flows are visualized by the positions of passive marker particles from the previous 10 periods of time. As expected for the single-frequency case in Figure 1(a), the flow resembles the classical, four vortices associated with steady streaming. However, for two-frequency motion, shown in Figures 1(b,c), the vortices are no longer symmetric and there is a net horizontal flow. The bottom panels show the cylinder’s horizontal position for each case. In Figure 1(b), the net flow is left-to-right – away from the cylinder, the green particles (most current period) are to the right of the blue particles (past period), indicating a net rightward flow. See movie 1 in the supplementary materials to compare the flow patterns of single-frequency, two-frequency, and reverse two-frequency oscillations. We note that because of the periodic boundary conditions these simulations correspond to a lattice of oscillating cylinders.
@@ 3.2. Pumping in a Channel
The results in 1(b) suggest that cylinders oscillating with two frequencies can be used to pump fluids. To test that idea more directly, we repeated the computational experiments in a channel of height 8 and length 32 with no-slip boundary conditions on the top and bottom and periodic conditions in the horizontal direction. Figure 2 is a picture of movie 2 at time $T=250$, and they show the contrast of single- and two-frequency motions inside a channel. Again, for the single-frequency case, the four vortices align symmetrically around the cylinder, and particles away from the cylinder remain stationary. However, when the cylinder oscillates with two-frequencies, the vortices lose their symmetry, and fluid is pumped to the right. Applying the negative waveform again reversed the direction of motion (results not shown).
@@
Figure 3(a) shows the flux through a channel of height $H$,
@@
over $5$ periods for both the single- and two-frequency cases for amplitude $\epsilon=0.9$. The flux for the two-frequency case shows a nonzero average flux on the scale of about 10% of the amplitude of the oscillation. We quantify pumping by the time average of the flux over the period $T$:
$\langle Q(t)\rangle\;=\frac{1}{T}\int_{t-T}^{t}Q(\tau)\;d\tau.$ (9)
@@
Figure 3(c) shows that the steady flux grows like amplitude cubed. For single-frequency oscillations the magnitude of the steady flow is second order in amplitude and the oscillatory component is first order (Holtsmark et al., 1954; Wang, 1968; Chong et al., 2013). In Figure 3(d) we show that for the two-frequency case these same scalings for the size of the steady and oscillatory flows result as measured by the $2$-norm of the period-averaged flow and the time average of the $2$-norm of the velocity, respectively. Previously, analysis on steady streaming has only been carried out to second order in amplitude, and so to derive the terms responsible for pumping, we need to extend the analysis to third order.
Figure 3. (a) Flux vs. time is presented for both the single- and two-frequency case for amplitude 0.9 and Reynolds number 10. (b) Time-averaged fluxes are shown for varying amplitudes $(\epsilon$). (c) The steady time-averaged flux is third order in amplitude for two-frequency oscillations. (d) The size of the oscillatory component, $\langle\|u\|_{2}\rangle$, is first order in amplitude, and the size of the steady streaming flow, $\|\langle u\rangle\|_{2}$, is second order in amplitude for the two-frequency oscillation.
3.3. Effect of Reynolds Number on Pumping
Figure 4. At amplitude $\epsilon=0.7$, frequency ratio $\alpha=2$, and time $\text{T}=100$, the time-averaged fluxes for a 32 by 8 channel are measured for various Reynolds numbers. Beyond the critical Reynolds number (Re $\approx 40$) at which a boundary layer forms, the time-averaged fluxes appear to scale as $\sqrt{\text{Re}}$.
We repeat the computational studies from the previous section to examine how the flux and flow structure change with Reynolds numbers. We note that at $\text{Re}=0$, the time-averaged flow is zero because of the dynamic reversibility of Stokes flow, and thus there is no pumping at zero Reynolds number. Figure 4 shows the time-averaged flux for $10\leq\text{Re}\leq 100$ at amplitude $\epsilon=0.7$, frequency ratio $\alpha=2$, and time $\text{T}=100$. The flux grows with Reynolds number and shows a change in scaling around $\text{Re}=40$, above which the time-averaged flux scales approximately as $\sqrt{\text{Re}}$. This change in scaling coincides with the Reynolds number at which a boundary develops in the single-frequency case (Chong et al., 2013).
Figure 5. For a 32 by 8 channel at amplitude $\epsilon=0.7$ and time $\text{T}=100$, the streamlines and the flow directions (normalized velocity field) of the time-averaged flow are shown at frequency ratios $\alpha=1$ (a, c) and $\alpha=2$ (b, d) and $\text{Re}=10$ (a, b) and $\text{Re}=50$ (c, d).
In Figure 5, we show the streamlines and normalized velocity vectors of the time-averaged flow for both single and two-frequency cases at Re = 10 and Re = 50 which are above and below the critical Reynolds number at which a boundary layer forms. At Re=10 for the single-frequency case we see the classical quadrupole-like structure where the red streamlines indicate counterclockwise flow, and the blue indicates clockwise flow (Figure 5(a)). For frequency ratio $\alpha=2$, there is a clear flow to the right, and vortices are no longer left-right symmetric. At Re = 50, there is a boundary layer present for both the single and two-frequency cases consisting of four rotating vortices with counter-rotating vortices outside the layer (Figure 5(c-d)). For two-frequency oscillations the net flow passes between the inner and outer vortices downstream from the cylinder.
@@ 4. Low Amplitude Analysis
We now analyze the problem in the limit of small amplitude $(\epsilon\ll 1)$, and we restrict our analysis to small streaming Reynolds numbers $(\text{Re}_{s}\ll 1)$ to apply a regular perturbation. The single-frequency solution has been previously computed through second order in amplitude (Holtsmark et al., 1954; Chong et al., 2013), but extending the analysis to third order is not analytically tractable. However, we derive the structure of the higher order solutions, and, from the structure alone, we can understand why pumping is a third order effect for the two-frequency case.
@@ 4.1. Governing Equations
We solve the problem in the reference frame in which the cylinder is stationary with an oscillatory flow at infinity. The boundary conditions in terms of the stream function are
@@ 4.2.1. Single-Frequency Oscillation
$\psi_{1}=\Re\left(a_{1}(r)e^{-it}\right)\sin(\theta),$ (24)
where $\Re\left(z\right)$ denotes the real part of $z$. Analytically, $a_{1}(r)$ is a solution composed of Hankel functions (Watson, 1922). We substitute (24) into the quadratic nonlinearity on the right side of Equation (18) to determine the form of the solution at second order. This computation involves all products of $e^{-it}$ and $e^{it}$, which results in a steady term and terms proportional to $e^{\pm i2t}$. Therefore, at second order, we obtain
@@ 4.2.2. Two-Frequency Oscillation
We express the solution in this second form with three pairs of terms grouped together to highlight the origin of the different terms in the sum. The first two pairs involving a steady term and unsteady term correspond to the single-frequency solutions from each of the two driving frequencies, and the last pair of unsteady terms arises from the interaction between the driving frequencies. Thus as has been reported previously, the steady flow at second order is the superposition of two streaming fields caused by the individual, single-frequency oscillations (Davidson and Riley, 1972; Kotas et al., 2008).
@@
Unlike the single-frequency solution (26), there are steady terms at third order in amplitude. In the following sections, we explain how the third order, steady terms produce the observed left-right asymmetry and are responsible for producing a net force that pumps the fluid.
@@ 4.3. Symmetry of Steady Flow
To illustrate how the steady flows at third order create asymmetry, consider the steady solution,
@@
We show that the familiar second order flow has a different spatial symmetry than the third order flow that arises from the interactions of the flows generated by the two different driving frequencies. Specifically, the spatial symmetry arises from the $\theta$-dependence of the stream functions. To illustrate these symmetries, we examine the streamlines of the three stream functions,
@@
This choice of $f(r)$ satisfies the boundary conditions on the cylinder and decays to zero at infinity, but it is not related to the solution of the equations.